Define T:R^3 -R^3 by
T(x1,x2,x3) = (x1+x2,x2+x3,x1-x3,2x1+x2-x3)
Check that T is a linear operator . Find the kernal and range of T. Find the dimension of the kernal.
1
Expert's answer
2020-04-27T18:31:45-0400

To show, that T is a linear operator let's check the following property:

"T(\lambda_1 v_1 + \lambda_2 v_2) = \lambda_1T(v_1) + \lambda_2 T(v_2)", where "v_1 = (x_1, y_1, z_1), v_2 = (x_2, y_2, z_2)". Indeed:

"T(\lambda_1 v_1 + \lambda_2 v_2) = T((\lambda_1 x_1, \lambda_1 y_1, \lambda_1 z_1)+ (\lambda_2 x_2, \lambda_2 y_2, \lambda_2 z_2))="

"= T(\lambda_1 x_1 + \lambda_2 x_2, \lambda_1 y_1 + \lambda_2 y_2, \lambda_1 z_1 + \lambda_2z_2)" "= (\lambda_1(x_1 + y_1) + \lambda_2(x_2 + y_2), \lambda_1(y_1 + z_1) + \lambda_2(y_2 + z_2), \lambda_1(x_1 -z_1) + \lambda_2(x_2 - z_2), \lambda_1(2x_1 + y_1 - z_1) + \lambda_2(2x_2 + y_2 - z_2)) ="

"= \lambda_1(x_1 + y_1, y_1 + z_1, x_1 - z_1, 2x_1 + y_1 - z_1) + \lambda_2((x_2 + y_2, y_2 + z_2, x_2 - z_2, 2x_2 + y_2 - z_2) = \lambda_1 T(v_1) + \lambda_2T(v_2)"

Let's now find the kernal and range of T.

"Ker(T) = v \in R^3 | T(v) = 0" , so:

"T(v) = T(x, y, z) = (x+y, y+z, x-z, 2x+y-z) = 0." This means, that:

"\begin{cases} 
x+y = 0 \\
y+z = 0 \\
x-z = 0\\
2x+y-z = 0
\end{cases}" Thus "\begin{cases}
x+y=0\\
z= x \\

\end{cases}" . Hence, "Ker (T) = Lin (1,-1,1), dim(Ker (T)) = 1."


Now let's do the same for the range:

"T(v) = T(x, y, z) = (x+y, y+z, x-z, 2x+y-z) ="

"= x(1, 0, 1, 2) + y(1,1,0,1)+z(0,1,-1,-1)." But, since "(0,1,-1,-1) = -(1,0,1,2) + (1,1,0,1)," we obtain that: "T(x,y,z) = (x-z)(1,0,1,2) + (y+z)(1,1,0,1)," and:

"Range(T) = Lin((1,0,1,2),(1,1,0,1))."